A reheat cycle with two stages of reheating is executed, with steam expanding initially from 90 bar and 530°C. The two reheater pressures are 10 bar and 0.6 bar, the steam leaves each reheater at 480°C. Condensation occurs at 0.03 bar. For an ideal cycle, Find the following: a. Heat supplied and efficiency b. For the engine, ignore the pressure drop through the reheaters, let the engine operate through the same states and compute Work engine and efficiency. c. What is the steam flow rate for an engine output of 20,000KW?


1
Expert's answer
2021-05-07T08:12:20-0400

Heat input in boiler


Qin = m (hi-hig) + in (hg-h₂) + (m-mpm₂) (his-hy) = 3.02 (3360-3-865:22) + 3-02 (3534-4-29317)


+ (3-01-0-532-0-0716) (2584-3-2167.2) = 10824-5 ks/ (kW).


Ideal Rankine ile Estruny Net work Cisentropie)


530+4-113-34


10824-5 20.479 (47.9%)


Actual Rankine Cycle efficiency actional mode wolk


20.4448


4815-18


(44-48%)


& Plant Capacity - Actual Work (Power) generatorettons


24815-1880-950-9 = 4116.99 kw


= 4.12 MW


(5) Heat released in condenser = = (3.02- 0.094 - 0.843) (2561.7- 251.12) 5442.2 KJ/s


Heat rejected by steam = Heat absorbed by boiling water = = mass × shc × ∆T


5442.2= mw × 4.187 × (21-10)

mass = 118.16 kg/s